# Open the Eyes of MPNN: Vision Enhances MPNN in Link Prediction

### v2

@@ 4.3 Compatible Features Integration
Algorithm 1 The GVN Framework (Attention-based)
@@ 4.4 E-GVN: An Efficient Variant
Considering the scalability of GVN on large-scale graphs, we propose a more efficient variant called Efficient GVN (E-GVN), whose architecture is presented in Fig. 1 and the algorithm is in Appendix A.2. The specialized modifications are introduced as follows.
@@
Partial-adaptive VSFs. GVN use end-to-end fintuning of vision encoder $\mathrm{VE}_{\psi}$ to refine the VSFs. However, such end-to-end fintuning requires retaining all visualization images and thus lead to substantial memory overhead. To enhance scalability in terms of space, E-GVN first freezes the vision encoder to extract static VSF $\mathbf{v}_{v}$. Instead, those static VSFs $\{\mathbf{v}_{v}\}$ can be stored and loaded as a fixed vector repository, significantly reducing the storage payload. After that, as we still want the VSFs to flexibly adapt to the different scenarios, we additionally append an trainable Adapter with learnable parameters $\phi$, which processes the static VSFs $\mathbf{v}_{v}$ to adaptive $\tilde{\mathbf{v}}_{v}$, where $\tilde{\mathbf{v}}_{v}=\text{Adaptor}_{\phi}(\mathbf{v}_{v})$ for each node $v$.
Pre-MPNN Integration on Attributes. Unlike GVN, which integrates VSFs after the MPNN, E-GVN incorporates $\tilde{\mathbf{v}}_{v}$ into the node attribute $\mathbf{x}_{v}$ to produce the vision-aware node attribute $\tilde{\mathbf{x}}_{v}$, which are then fed into the MPNN as inputs. Due to the integration occurs before message passing, this approach allows the vision-aware structural information to participate in the message propagation and aggregation to get further enhanced.
@@
Time Complexity. Let $n$ and $l$ denote the numbers of nodes and links, respectively, $F$ and $F^{\prime}$ be the dimensions of node features and MPNN-produced representations, respectively, $O(\text{MPNN})$ and $O(\text{VE})$ be the complexities of the MPNN and vision encoder, respectively, and $S$ be the dimension of the VSF. With default attention-based integration, the time complexity of GVN and E-GVN are $O(\text{MPNN})+lO(\text{VE})+O(l(SF^{\prime}+F^{\prime 2}))$ and $O(\text{MPNN})+nO(\text{VE})+O(n(S^{2}+SF+F^{2}))$, respectively. Here we can see that E-GVN reduces the computational complexity, as $n$ is much less than $l$. More details are provided in Appendix B.
5 Experiment
@@ 5.1 Evaluation on Real-World Datasets
Datasets. We conduct experiments on seven widely used link prediction benchmarking, including Planetoid citation networks: Cora (McCallum et al., 2000), Citeseer (Sen et al., 2008), and Pubmed (Namata et al., 2012), and the large-scale OGB link prediction datasets (Hu et al., 2020): ogbl-collab, ogbl-ppa, ogbl-citation2 and ogbl-ddi. Statistics of those datasets are shown in Appendix C.
@@
Baselines. Baseline methods used include 1) Link prediction heuristics: Common Neighbor counts (CN) (Barabási & Albert, 1999), Adamic-Adar (AA) (Adamic & Adar, 2003), and Resource Allocation (RA) (Zhou et al., 2009); 2) MPNNs: GraphSAGE (Hamilton et al., 2017) and Graph Convolutional Network (GCN) (Kipf & Welling, 2016); 3) Advanced SF-enhanced MPNNs: SEAL (Zhang & Chen, 2018a) and NBFNet (Zhu et al., 2021) with path-based SFs, Neo-GNN (Yun et al., 2021), BUDDY (Chamberlain et al., 2023), and NCNC (Wang et al., 2024) that is the SOTA SF-enhanced MPNN.
@@
Performance Evaluation. The use of evaluation metrics follows (Chamberlain et al., 2023; Wang et al., 2024). Specifically, for the Planetoid datasets, we use the hit-ratio at 100 (HR@100), while for the OGB datasets, we use the metrics in their official documents, [^3] https://ogb.stanford.edu/docs/leader_linkprop/. i.e., hit-ratio at 50 (HR@50) for ogbl-collab, HR@100 for ogbl-ppa, Mean Reciprocal Rank (MRR) for ogbl-citation2, and hit-ratio at 20 (HR@20) for ogbl-ddi. [^4] Evaluation results on other metrics are in Appendix F. All results are reported as average and standard variation over 10 trials with different random seeds.
@@
Implementation Details. All experiments are conducted on an NVIDIA A100 80G GPU. For GVN and E-GVN, the candidate hyperparameters include the visual perception scope (i.e., the subgraph visualization hop) $k$ ranging from 1 to 3, the hidden dimension ranging from 512 to 2048, the number of MPNN layers and readout predictor layers varying from 1 to 3, the two separate learning rates for trainable VSFs and compatible feature integration among {$10^{-7}$, $10^{-5}$, 0.001, 0.01}, and the weight decay from 0 to 0.0001. The hyperparameters with the best validation accuracy are selected. For the model parameters, we utilize the Adam algorithm (Kingma, 2014) as the optimizer. As default settings, we use Graphviz (Gansner & North, 2000) as the graph visualizer and use a pretrained ResNet50 (He et al., 2016) as the vision encoder. More details on the experimental setup are put in Appendix D.
@@
Configurations of the Proposed Methods. We use four configurations of the GVN and E-GVN frameworks: GVNGCN, GVNNCNC, E-GVNGCN, and E-GVNNCNC, as shown in Table 3. The subscript (i.e., ‘GCN’ or ‘NCNC’) indicates the MPNN model used within the frameworks. Our choice of MPNNs is based on the following considerations: First, using GCN, a classic and straightforward MPNN, within the proposed frameworks directly demonstrates the enhancement provided by vision awareness when compared to MPNN alone. This allows us to clearly see the impact of incorporating the visual modality into MPNNs. Second, based on NCNC, a state-of-the-art MPNN enhanced by advanced structural features, the comparison between GVNNCNC/E-GVNNCNC and NCNC will highlight the additional performance improvements from the vision awareness that are orthogonal to existing advanced methods. This demonstrates how visual features can complement and enhance existing sophisticated models. Although we focus on those two MPNNs, the proposed GVN and E-GVN are compatible with most MPNN models.
@@
(i) By integrating visual structural features into GCN, GVN enhances GCN with relative improvements (HR@100) on link prediction performance by 21.47%, 25.13%, and 38.00% on Cora, Citeseer and PubMed, respectively. Similarly, the E-GVN framework achieved relative improvements of 19.79%, 23.51%, and 35.68% upon GCN on those three datasets. Additionally, due to the efficient designs, E-GVN extends the visual benefits to large-scale graph datasets ogbl-collab, ogbl-ppa, ogbl-citation2 and ogbl-ddi, bringing respective 38.86% (HR@50), 72.20% (HR@100), 1.60% (MRR), and 62.42% (HR@20) relative improvement. Therefore, consistent results indicate that visual awareness has been effectively integrated into GCN by the proposed methods and significantly enhances the link prediction performance of GCN.
@@
(ii) Based on SF-enhanced NCNC model, both GVNNCNC and E-GVNNCNC still show remarkable improvements. Although NCNC has been a leading link prediction model by enhancing MPNN with dedicated common neighbor structural features, adopting the GVN and E-GVN frameworks brings additional improvements over NCNC. These results demonstrate that the visual awareness provided by the GVN or E-GVN frameworks own advantages not covered by existing methods and can be used in conjunction with current SOTA methods for better performance. Therefore, GVNNCNC and E-GVNNCNC surpass all baselines and are established as more powerful link prediction approaches.
@@
(iii) Compared to GVN, the computationally efficient design of E-GVN makes it more suitable for large-scale graphs. GVNGCN is more powerful than E-GVNGCN. This may be because the visual structural awareness in GVN is link-based, therefore explicitly revealing link pairwise relations such as common neighbors or shortest path differences for distinguishing links with isomorphic nodes better. However, when using SF-enhanced NCNC, the improvements brought by E-GVNNCNC surpass those of GVNNCNC. We attribute this to the fact that, in GVNNCNC, both the link-centered visual awareness and the common-neighbor-based SF can reveal the common neighbor information between the link nodes, which makes their effects partially overlap. In contrast, the pre-MPNN integration in E-GVN allows vision-aware structural information to be further aggregated during message passing, which provides more comprehensive visual awareness.
5.2 Scalability Analysis
Fig. 5 compares the inference time and GPU memory for inferring one batch of samples from Cora (including pre-processing time). Among all the methods in comparison, GVN is the most time-consuming, followed by SEAL and NBFnet. Those three methods also require considerably more memory than the others. This elevated resource consumption is due to the need for pre-processing and computation for each link, and the storage of intermediate variables with respect to links. Particularly, GVN requires graph visualization, which introduces extra pre-processing time. Therefore, similar to SEAL and NBFnet, GVN is not well-suited for large-scale graph computations. In contrast, the visualization of E-GVN is node-dependent, which can be reused for diverse links. After amortizing the pre-processing visualization time across link queries, E-GVN still exhibits computational overhead similar to the base models GCN and NCNC. Therefore, when leveraging the lightweight base models (e.g., GCN or NCNC), E-GVN maintains efficiency for large-scale graphs.
@@ 5.3 Ablation and Sensitivity Analysis
Visualization Style Consistency. Table 4 shows that inconsistent styles degrade performance and emphasize keeping styles consistent. The inconsistent styles are sampled from a fixed range of node colors, shapes, and graph visualizers, as detailed in Appendix H. In Appendix I, we have included more constructive findings by studying the impacts of node colors (Appendix I.1), node shapes (Appendix I.2), and node labeling strategies (Appendix I.3).
@@
Feature Integration Strategies. Table 8 shows the performance when varying feature integration strategies in Section 4.3. Experimental results suggest that, in most cases, the default attention-based integration (denoted by “Attention” in Table 8) is more effective than concatenated integration (denoted by “Concat”) and weighted integration (denoted by “Weight”).
VSF Adaptivity. Table 9 shows the impacts of VSF Adaptivity, where “freeze” means using a frozen vision encoder, making VSFs static, “partial” adopts the partial adaptive VSFs proposed in Section 4.4 that freezes the vision encoder but supplements it with a learnable adapter to provide some adaptivity, and “full” means the models are fully fine-tuned. Results show that static VSFs lose the adaptivity, leading to performance degradation. The “partial” strategy can provide comparable performance with the “full” strategy, demonstrating its effectiveness.
@@ 5.4 Comparison with Positional Encodings
Except comparing with heuristic SFs, to further exhibit the effectiveness of VSFs, we here compared VSFs for nodes with 4 representative encoding mechanisms: 1) 2-d axis position of nodes in subgraph image, 2) Laplacian positional encoding, 3) Distances to other nodes, and 4) Node degree (centrality encoding). These positional encodings (PE) were used as node features and decoded by a 2-layer GCN with a 2-layer MLP for link prediction. The results are presented in Table 10. As can be seen, VSFs outperform the PEs across datasets. Notably, directly encoding the 2-dimensional axis is not effective enough, underscoring the importance of using a vision encoder for comprehensive structural information.
@@ Acknowledgement
This work was supported by National Key Research and Development Program of China (No. 2022YFA1004102), NSFC grant (No. 62136005), and the Research Grants Council of the Hong Kong Special Administrative Region (Grants 16200021, 16202523, C7004-22G-1). We also acknowledge the authors of (Chen & Wei, 2025; Lu et al., 2023, 2024, 2025).

### v3

@@ 4.3 Compatible Features Integration
Algorithm 1 The GVN Framework (Attention-based).
@@ 4.4 E-GVN: An Efficient Variant
Considering the scalability of GVN on large-scale graphs, we propose a more efficient variant called Efficient GVN (E-GVN), whose architecture is presented in Figure 1 and the algorithm is in Appendix A.2. The specialized modifications are introduced as follows.
@@
Partial-adaptive VSFs. GVN use end-to-end finetuning of vision encoder $\mathrm{VE}_{\psi}$ to refine the VSFs. However, such end-to-end finetuning requires retaining all visualization images and thus lead to substantial memory overhead. To enhance scalability in terms of space, E-GVN first freezes the vision encoder to extract static VSF $\mathbf{v}_{v}$. Instead, those static VSFs $\{\mathbf{v}_{v}\}$ can be stored and loaded as a fixed vector repository, significantly reducing the storage payload. After that, as we still want the VSFs to flexibly adapt to the different scenarios, we additionally append a trainable Adapter with learnable parameter $\phi$, which processes the static VSF $\mathbf{v}_{v}$ to adaptive $\tilde{\mathbf{v}}_{v}$, where $\tilde{\mathbf{v}}_{v}=\text{Adaptor}_{\phi}(\mathbf{v}_{v})$ for each node $v$.
Pre-MPNN Integration on Attributes. Unlike GVN, which integrates VSFs after the MPNN, E-GVN incorporates $\tilde{\mathbf{v}}_{v}$ into the node attribute $\mathbf{x}_{v}$ to produce vision-aware node attribute $\tilde{\mathbf{x}}_{v}$, which are then fed into the MPNN as inputs. Due to the integration occurs before message passing, this allows the vision-aware structural information to participate in the message propagation and aggregation to get further enhanced.
@@
Time Complexity. Let $n$ and $l$ be the numbers of nodes and links, respectively, $F$ and $F^{\prime}$ be the dimensions of node features and MPNN-produced representations, respectively, $O(\text{MPNN})$ and $O(\text{VE})$ be the complexities of the MPNN and vision encoder, respectively, and $S$ be the dimension of the VSF. With the default attention-based integration, the time complexities of GVN and E-GVN are $O(\text{MPNN})+lO(\text{VE})+O(l(SF^{\prime}+F^{\prime 2}))$ and $O(\text{MPNN})+nO(\text{VE})+O(n(S^{2}+SF+F^{2}))$, respectively. We can see that E-GVN reduces the computational complexity, as $n$ is much less than $l$. More details are provided in Appendix B.
5 Experiments
@@ 5.1 Evaluation on Real-World Datasets
Datasets. We conduct experiments on seven widely used link prediction benchmarking, including the Planetoid citation networks: Cora (McCallum et al., 2000), Citeseer (Sen et al., 2008), and Pubmed (Namata et al., 2012), and the large-scale OGB link prediction datasets (Hu et al., 2020): ogbl-collab, ogbl-ppa, ogbl-citation2 and ogbl-ddi. Statistics of these datasets are shown in Appendix C.
@@
Baselines. Baseline methods used include 1) Link prediction heuristics: Common Neighbor counts (CN) (Barabási & Albert, 1999), Adamic-Adar (AA) (Adamic & Adar, 2003), and Resource Allocation (RA) (Zhou et al., 2009); 2) MPNNs: GraphSAGE (Hamilton et al., 2017) and Graph Convolutional Network (GCN) (Kipf & Welling, 2017); 3) Advanced SF-enhanced MPNNs: SEAL (Zhang & Chen, 2018) and NBFNet (Zhu et al., 2021) with path-based SFs, Neo-GNN (Yun et al., 2021), BUDDY (Chamberlain et al., 2023), and NCNC (Wang et al., 2024) that is the SOTA SF-enhanced MPNN.
@@
Performance Evaluation. The use of evaluation metrics follows (Chamberlain et al., 2023; Wang et al., 2024). Specifically, for the Planetoid datasets, we use the hit-ratio at 100 (HR@100), while for the OGB datasets, we use the metrics in their official documents, [^4] https://ogb.stanford.edu/docs/leader_linkprop/. i.e., hit-ratio at 50 (HR@50) for ogbl-collab, HR@100 for ogbl-ppa, Mean Reciprocal Rank (MRR) for ogbl-citation2, and hit-ratio at 20 (HR@20) for ogbl-ddi. [^5] Evaluation results on other metrics are in Section 5.2. All results are reported as average and standard variation over 10 trials with different random seeds.
@@
Implementation Details. All experiments are conducted on an NVIDIA A100 80G GPU. For GVN and E-GVN, the candidate hyperparameters include the visual perception scope (i.e., the subgraph visualization hop) $k$ ranging from 1 to 3, the hidden dimension ranging from 512 to 2048, the number of MPNN layers and readout predictor layers varying from 1 to 3, the two separate learning rates for trainable VSFs and compatible feature integration among {$10^{-7}$, $10^{-5}$, 0.001, 0.01}, and the weight decay from 0 to 0.0001. The hyperparameters with the best validation accuracy are selected. For the model parameters, we utilize the Adam algorithm (Kingma, 2014) as the optimizer. As default settings, we use Graphviz (Gansner & North, 2000) as the graph visualizer and use a pretrained ResNet50 (He et al., 2016) as the vision encoder. More details on the experimental setup are in Appendix D.
Table 4: The performance comparison of the proposed methods against the strongest baseline NCNC across broader metrics.
Cora Citeseer Pubmed Collab PPA Citation2 DDI
hit@3 GVNNCNC $26.66_{\pm 5.96}$ $59.97_{\pm 6.21}$ $\mathbf{32.23_{\pm 5.69}}$
hit@10 GVNNCNC $\mathbf{58.83_{\pm 5.29}}$ $75.28_{\pm 3.03}$ $40.34_{\pm 2.28}$
hit@20 GVNNCNC $\mathbf{70.01_{\pm 4.44}}$ $81.11_{\pm 1.30}$ $53.33_{\pm 2.67}$
hit@50 GVNNCNC $82.06_{\pm 1.94}$ $88.88_{\pm 0.98}$ $\mathbf{71.66_{\pm 2.75}}$
hit@100 GVNNCNC $90.70_{\pm 0.56}$ $94.12_{\pm 0.58}$ $82.17_{\pm 0.77}$
mrr GVNNCNC $\mathbf{24.66_{\pm 4.51}}$ $62.74_{\pm 6.63}$ $26.32_{\pm 6.67}$
@@
Configurations of the Proposed Methods. We use four configurations of the GVN and E-GVN frameworks: GVNGCN, GVNNCNC, E-GVNGCN, and E-GVNNCNC, as shown in Table 3. The subscript (i.e., ‘GCN’ or ‘NCNC’) indicates the MPNN model used within the frameworks. Our choice of MPNNs is based on the following considerations: First, using GCN, a classic and straightforward MPNN, within the proposed frameworks directly demonstrates the enhancement provided by vision awareness when compared to MPNN alone. This allows us to clearly see the impact of incorporating the visual modality into MPNNs. Second, based on NCNC, a state-of-the-art MPNN enhanced by advanced structural features, the comparison between GVNNCNC/E-GVNNCNC and NCNC will highlight the additional performance improvements from the vision awareness that are orthogonal to existing advanced methods. This demonstrates how visual features can complement and enhance existing sophisticated models. Although we focus on these two MPNNs, the proposed GVN and E-GVN are compatible with most MPNN models.
@@
(i) By integrating visual structural features into GCN, GVN enhances GCN with relative improvements on link prediction performance (HR@100) by 21.47%, 25.13%, and 38.00% on Cora, Citeseer and PubMed, respectively. Similarly, E-GVN achieves relative improvements of 19.79%, 23.51%, and 35.68% upon GCN on these three datasets. Additionally, due to the efficient designs, E-GVN extends the visual benefits to large-scale graph datasets ogbl-collab, ogbl-ppa, ogbl-citation2 and ogbl-ddi, bringing relative improvements of 38.86% (HR@50), 72.20% (HR@100), 1.60% (MRR), and 62.42% (HR@20), respectively. Therefore, consistent results indicate that visual awareness has been effectively integrated into GCN by the proposed methods and significantly enhances the link prediction performance of GCN.
@@
(ii) Based on SF-enhanced NCNC model, both GVNNCNC and E-GVNNCNC still show remarkable improvements. Although NCNC is a leading link prediction model by enhancing MPNN with dedicated common neighbor structural features, adopting GVN and E-GVN brings additional improvements over NCNC. These results demonstrate that the visual awareness provided by GVN or E-GVN own advantages not covered by existing methods and can be used in conjunction with current SOTA methods for better performance. Therefore, GVNNCNC and E-GVNNCNC surpass all baselines and are established as more powerful link prediction approaches.
@@
(iii) Compared to GVN, the computationally efficient design of E-GVN makes it more suitable for large-scale graphs. GVNGCN is more powerful than E-GVNGCN. This may be because the visual structural awareness in GVN is link-based, therefore explicitly revealing link pairwise relations such as common neighbors or shortest path differences for distinguishing links with isomorphic nodes better. However, when using SF-enhanced NCNC, the improvements brought by E-GVNNCNC surpass those of GVNNCNC. We attribute this to the fact that in GVNNCNC, both the link-centered visual awareness and the common-neighbor-based SF can reveal the common neighbor information between the link nodes, which makes their effects partially overlap. In contrast, the pre-MPNN integration in E-GVN allows vision-aware structural information to be further aggregated during message passing, which provides more comprehensive visual awareness.
5.2 Evaluation on Broader Metrics
Besides the metrics listed in Table 3, the performances of proposed methods against the strongest baseline NCNC with broader metrics are presented in Table 4. In total, E-GVN achieves 39 best scores (in bold), GVN achieves 7 best scores, and our strongest baseline NCNC achieves 3 best scores. Therefore, GVNNCNC and E-GVNNCNC consistently outperform NCNC in more comprehensive metrics.
@@ 5.3 Ablation and Sensitivity Analysis
Visualization Style Consistency. Table 5 shows that inconsistent styles degrade performance and emphasize keeping styles consistent. The inconsistent styles are sampled from a fixed range of node colors, shapes, and graph visualizers, as detailed in Appendix G. In Appendix H, we include more comprehensive findings by studying the impacts of node colors (Appendix H.1), node shapes (Appendix H.2), and node labeling strategies (Appendix H.3).
@@
Feature Integration Strategies. Table 9 shows the performance when varying feature integration strategies in Section 4.3. Experimental results suggest that in most cases, the default attention-based integration (denoted “Attention” in Table 9) is more effective than concatenated integration (denoted “Concat”) and weighted integration (denoted “Weight”).
VSF Adaptivity. Table 10 shows the impacts of VSF Adaptivity, where “freeze” means using a frozen vision encoder, making VSFs static, “partial” adopts the partial adaptive VSFs proposed in Section 4.4 that freezes the vision encoder but supplements it with a learnable adapter to provide some adaptivity, and “full” means the models are fully fine-tuned. Results show that static VSFs lose adaptivity, leading to performance degradation. The “partial” strategy can provide comparable performance with the “full” strategy, demonstrating its effectiveness.
@@
5.4 Scalability Analysis
Figure 5 compares the inference time and GPU memory for inferring one batch of samples from Cora (including pre-processing time). Among all the methods in comparison, GVN is the most time-consuming, followed by SEAL and NBFnet. These three methods also require considerably more memory than the others. This elevated resource consumption is due to the need for pre-processing and computation for each link, and the storage of intermediate variables with respect to links. Particularly, GVN requires graph visualization, which introduces extra pre-processing time. Therefore, similar to SEAL and NBFnet, GVN is not well-suited for large-scale graph computations. In contrast, the visualization of E-GVN is node-dependent, which can be reused for diverse links. After amortizing the pre-processing visualization time across link queries, E-GVN still exhibits computational overhead similar to the base models GCN and NCNC. Therefore, when leveraging the lightweight base models (e.g., GCN or NCNC), E-GVN maintains efficiency for large-scale graphs.
@@ 5.5 Comparison with Positional Encodings
In addition to comparing with heuristic SFs, we further demonstrate the effectiveness of VSFs by comparing them against four representative encoding mechanisms for nodes: 1) 2-dimensional coordinates of nodes in the subgraph image, 2) Laplacian positional encoding, 3) Distances to other nodes, and 4) Node degree (centrality encoding). These positional encodings (PEs) are utilized as node features and processed by a 2-layer GCN followed by a 2-layer MLP for link prediction. The comparison results of Hits@100 score among VSFs and these encoding mechanisms are presented in Table 11. As observed, VSFs outperform the PEs across various datasets. Notably, directly encoding the 2-dimensional coordinates proves insufficient, highlighting the importance of employing a vision encoder to capture comprehensive structural information.
@@ Acknowledgement
This work was supported by National Key Research and Development Program of China (No. 2022YFA1004102), NSFC grant (No. 62136005), and the Research Grants Council of the Hong Kong Special Administrative Region (Grants 16200021, 16202523, C7004-22G-1). We also acknowledge the authors of (Chen & Wei, 2025; Lu et al., ; Lu et al., 2024, 2025).